OpenStudy (anonymous):

y varies inversely as twice x. When x = 1, y = 3. Find y when x = 4.

OpenStudy (rulnick):

y=k(1/(2x)) 3=k(1/2(1))=k(1/2) k=6 so y=6/(2x)=3/x=3/4

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

When x = 1, y = 3. Find y when x = 4. y = k/2x -> k = 2x y = 2 * 3 = 6 => y = 6/ 2 * 4 = 3/4